A positive charge of 5.00 μC is fixed in place. From a distance of 4.60 cm a from the fixed positive charge, a particle of mass 6.40 g and charge +3.60 μC is released from rest. What is the speed of the +3.60 μC when it is 9.00 cm away from the fixed positive charge?
Answer #1

using conservation of energy

electric potential energy = kinetic energy

k q1 q2 * ( 1/ r1 - 1/r2) = 0.5 m v^2

9* 10^9* 5* 3.6* 10^-12* ( 1/ 0.046 - 1/ 0.09) = 0.5* 6.4* 10^-3 v^2

v = 23.196 m/s

=========

Comment before rate in case any doubt, will reply for sure.. goodluck